I think I got this one figured out but again a pair of expert eyes would be great and extremely appreciated. Thanks!! Andrea
If I invest $1000 on May 18th in a savings account that pays 10% interest compounded every month, and I invest $1000 (on the same day) in an account which pays 10% interest that is compounded continuously:
(a)
Which will be worth more in one year?
Monthly
A = P(1 + r)^t
A = 1000(1 + .1)^12
A = $3138.43
Continuously
A = Pe^(rt)
A = 1000e^((.1)(1))
A = $1105.17
(b)
By how much?
$3138.43 – $1105.17 = $2033.26

Answer by venugopalramana(3286)

SEE MY COMMENTS BELOW
I think I got this one figured out but again a pair of expert eyes would be great and extremely appreciated. Thanks!! Andrea
If I invest $1000 on May 18th in a savings account that pays 10% interest compounded every month, and I invest $1000 (on the same day) in an account which pays 10% interest that is compounded continuously:
(a)
Which will be worth more in one year?
Monthly
A = P(1 + r)^t............NO....SEE BELOW
FORMULA IS A=P(1+R/N)^(NT)..WHERE N IS THE NUMBER OF PERIODS PER YEAR OVER WHICH COMPOUNDING IS DONE..HERE N=12..SINCE COMPOUNDING IS DONE EVERY MONTH..
SO WE GET A=1000(1+0.1/12)^12=1104.71
-----------------------------------------------------------
A = 1000(1 + .1)^12
A = $3138.43
Continuously
A = Pe^(rt)
A = 1000e^((.1)(1))
A = $1105.17...........OK ...CORRECT
B IS BETTER
By how much
$1105.17-1104.71 = $0.46 $=46 CENTS